OpenStudy (anonymous):

The hypotenuse of a right triangle has one end at the origin and one end on the curve y = x^5*e^(−5x), with x ≥ 0. One of the other two sides is on the x-axis, the other side is parallel to the y-axis. a) Express the area of the triangle in terms of x. b) Find the critical point for x>0. c) Find the maximum area of the triangle. Give an exact answer.

OpenStudy (anonymous):

The curve is: \[y=x^{5}e^{-5x}\]

OpenStudy (anonymous):

For part a)\[A=\frac{ 1 }{ 2 } (x*x^{5}e^{-5x})\]

OpenStudy (anonymous):

I am stuck on how to proceed with part b.

OpenStudy (anonymous):

First derivative test: first find what values of \(x\) make \[\dfrac{d}{dx}[x^5~e^{-5x}]=0\]

OpenStudy (anonymous):

\[\frac{ d }{ dx }(x^{5}e^{-5x})=e^{-5x}(5x^4-5x^5)\]

OpenStudy (anonymous):

I was getting x=1

OpenStudy (anonymous):

\(e^{-5x}>0\) for all \(x\), so solving \(5x^4-5x^5=0\) indeed gives you \(x=0,1\) (disregarding 0). Now check to see if \(x=1\) is a critical point. The sign of the derivative to either side of the point determines whether the curve is increasing or decreasing: On the interval \((0,1)\), \(y'>0\), so \(y\) is increasing on this interval. On \((1,\infty)\), \(y'<0\), so \(y\) is decreasing on this interval. The increase-decrease indicates a maximum at \(x=1\). And since \(y\) is continuous at \(x=1\), that makes \(x=1\) a critical point.

OpenStudy (anonymous):

Okay, thank you! The website kept saying x=1 was wrong so I'll be sure to contact my instructor about that.

OpenStudy (anonymous):

Maybe the answer is the point itself, which would be \((1,y(1))\) ?

OpenStudy (anonymous):

The box has it marked as x= and then I am supposed to type in the answer. I had wondered if I had to type the entire point but it says it only wants the x-value. Oh well. Thank you again!

OpenStudy (anonymous):

You're welcome